For this problem i just need to know the last equation.

A minor league baseball park has 7000 seats. Box seats cost $6,grandstand seats cost $4, and bleachers seats cost $2. When all seats are sold, the revenue is $26400. If the number of box seats is one-third the number of bleachers seats, how many seats of each type are there?

My 2 equations are

x + y + z = 7000
6x + 4y +2z = 26400

What is the last equation?

Found 3 solutions by solver91311, Edwin McCravy, stanbon:
Answer by solver91311(24713)



The number of box seats is one-third the number of bleacher seats, so:



Write back with your solution and I'll check it for you.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Edwin McCravy(20056)

For this problem i just need to know the last equation.

A minor league baseball park has 7000 seats. Box seats cost $6,grandstand seats cost $4, and bleachers seats cost $2. When all seats are sold, the revenue is $26400. If the number of box seats is one-third the number of bleachers seats, how many seats of each type are there?

My 2 equations are

x + y + z = 7000
6x + 4y +2z = 26400
What is the last equation?
It comes from this sentence:

"the number of box seats, x, is one-third the number 
of bleachers seats, z"

That says:

x is one-third of z 

or

x = ⅓z

and then you can clear of fractions by multiplying
both sides by 3 and getting 

3x = z

And you could rearrange it as

3x - z = 0

Then your system is

 x +  y + z =  7000
6x + 4y +2z = 26400
3x      - z =     0


answer:  x = 400, y = 5400, z = 1200

Edwin


Answer by stanbon(75887)

For this problem i just need to know the last equation.
A minor league baseball park has 7000 seats. Box seats cost $6,grandstand seats cost $4, and bleachers seats cost $2. When all seats are sold, the revenue is $26400. If the number of box seats is one-third the number of bleachers seats, how many seats of each type are there?
---
What is the last equation?
Seat Equation: x + y + z = 7000
Value Equation:6x + 4y +2z = 26400
Quantity Eq:: x = (1/3)z
==================================
Cheers,
Stan H.
=============